A new cruise ship line has just launched 3 new ahips; the Pacific Paradise, the Caribbean Paradise, and the Mediterranean Paradise. The Caribbean Paradise has 13 more deluxe staterooms than the Pacific Paradise. The Mesitterean Paradise has 40 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 615
Answer by JulietG(1812)

Let P be the Pacific
Let C be the Caribbean
Let M be the Mediterranean
.
Here is the information given in the problem:
C = P + 13
M = 4P - 40
C + P + M = 615
.
Let's rewrite the problem substituting the values given in the question.
C + P + M = 615
(P+13) + P + (4P-40) = 615
Add the Ps
6P + 13 - 40 = 615
Do the addition and subtraction
6P - 27 = 615
Add 27 to each side
6P = 642
Divide each side by 6
P = 107
.
If P is 107, then C is 107+13, or 120
If P is 107, then M is (4*107)-40, or 428-40, or 388
Add them together to make certain the problem is correct.
107 + 120 + 388 does indeed equal 615.
Success!
.
And now I'm ready to sail on the Princess Ruby!